# DNABERT-S: Learning Species-Aware DNA Embedding with Genome Foundation Models

### v1

@@ Abstract
Effective DNA embedding remains crucial in genomic analysis, particularly in scenarios lacking labeled data for model fine-tuning, despite the significant advancements in genome foundation models. A prime example is metagenomics binning, a critical process in microbiome research that aims to group DNA sequences by their species from a complex mixture of DNA sequences derived from potentially thousands of distinct, often uncharacterized species. To fill the lack of effective DNA embedding models, we introduce DNABERT-S, a genome foundation model that specializes in creating species-aware DNA embeddings. To encourage effective embeddings to error-prone long-read DNA sequences, we introduce Manifold Instance Mixup (MI-Mix), a contrastive objective that mixes the hidden representations of DNA sequences at randomly selected layers and trains the model to recognize and differentiate these mixed proportions at the output layer. We further enhance it with the proposed Curriculum Contrastive Learning (C2LR) strategy. Empirical results on 18 diverse datasets showed DNABERT-S’s remarkable performance. It outperforms the top baseline’s performance in 10-shot species classification with just a 2-shot training while doubling the Adjusted Rand Index (ARI) in species clustering and substantially increasing the number of correctly identified species in metagenomics binning. The code, data, and pre-trained model are publicly available at https://github.com/Zhihan1996/DNABERT_S.
@@ 1 Introduction
The introduction of genome foundation models, including DNABERTs and HyenaDNA [Ji et al., 2021; Dalla-Torre et al., 2023; Nguyen et al., 2023; Zhou et al., 2023], marks a transformative era in genomic analysis. These models, enriched with extensive domain knowledge through large-scale pre-training, have demonstrated impressive capabilities in various genome analysis tasks, especially where labeled data is available. However, a critical challenge persists: the lack of labeled data in key applications, such as metagenomics binning.
Metagenomics binning [Kang et al., 2015, 2019; Nissen et al., 2021; Meyer et al., 2022; Lamurias et al., 2023] aims to group DNA sequences from a complex mixture of multiple organisms, typically in the absence of prior labeling, which is essential for supervised fine-tuning. In this context, high-quality and discriminative DNA embeddings become an indispensable tool. Similarly, other important problems such as nucleosome positioning [Han et al., 2022] and DNA sequence comparison [Ren et al., 2022], and various biological analysis problems involving multi-modal data (e.g., gene, protein, pathway, etc.) [Ma et al., 2023] where direct model fine-tuning is not viable, also benefit substantially from effective DNA embeddings.
@@
Despite the critical role of DNA embeddings in genomic research, the current landscape reveals a notable shortfall in effective methods for their development. Current approaches primarily achieve DNA embeddings via three methods: 1) Descriptive textual features [Kang et al., 2015, 2019; Nissen et al., 2021], 2) Pre-trained Kmer embeddings [Ng, 2017; Ren et al., 2022; Han et al., 2022], and 3) Genome foundation models [Ji et al., 2021; Nguyen et al., 2023; Zhou et al., 2023]. The first two methods, while straightforward, fail to grasp complex semantic relationships. Moreover, existing genome foundation models, predominantly pre-trained with language modeling objectives [Radford et al., 2019; Devlin et al., 2018], often fail to yield representative embeddings [Reimers and Gurevych, 2019; Gao et al., 2021]. Our empirical analysis reveals that, in many scenarios, existing genome foundation models even underperform descriptive textual features, as detailed in Table 1.
@@
To address this challenge, we introduce DNABERT-S, a specialized genome foundation model tailored for generating species-aware DNA embeddings. As depicted in Figure 1, DNABERT-S distinguishes itself from other methods by its ability to effectively cluster and separate different species within the embedding space. This enhanced performance stems from the proposed Manifold Instance Mixup (MI-Mix) loss and Curriculum Contrastive Learning (C2LR) strategy. Contrastive learning enables the model to discern between similar and dissimilar DNA sequences, and curriculum learning incrementally presents more challenging training samples, fostering better learning and generalization. The training of DNABERT-S includes two phases: 1) differentiate similar and dissimilar sequences; 2) form more challenging contrastive anchors and continue the differentiation. In the first phase, we adopt a Weighted SimCLR Chen et al. [2020]; Zhou et al. [2022] training objective to encourage the model to group similar DNA sequences and separate dissimilar DNA sequences. In the second phase, we introduce Manifold Instance Mixup (MI-Mix) which mixes anchor instances at a randomly selected layer to create more challenging anchors for contrastive training.
@@
To evaluate DNABERT-S against existing methods, we compiled a comprehensive benchmark that included more than 1,000 species, reflecting the diversity of natural microbial communities. This benchmark includes complex datasets from CAMI2 [Meyer et al., 2022], a leading metagenomics binning benchmark, and extensive reference genomes from Genbank [Benson et al., 2012]
@@
We evaluate the methods on three types of tasks: metagenomics binning, clustering, and classification, from entirely unsupervised problems to few-shot learning. Experimental results indicate the remarkable performance of DNABERT-S over all the existing methods. Compared to the strongest existing method, DNABERT-S doubles its performance in the clustering task and achieves better performance with only $20\%$ of labeled data in the classification task (e.g., 2-shot v.s. 10-shot). Besides, in metagenomics binning, we show DNABERT-S is able to recover over $40\%$ and $80\%$ of species with an F1 score of over $0.5$ from respectively from synthetics and more realistic datasets, which is also one time more than the strongest baseline.
@@
Our contribution can be summarized as follows: 1) For the first time, we demonstrate the superiority of genome foundation models in learning effective DNA embeddings, opening new avenues for tackling a wide range of genomic research challenges; 2) We introduce DNABERT-S, a genome foundation model distinctly outperforming existing methods in learning DNA embeddings; 3) We introduce the Curriculum Contrastive Learning (C2LR) strategy with the Manifold Instance Mixup (MI-Mix) loss, which effectively facilitate DNA embedding learning; 4) We construct a large-scale evaluation benchmark for DNA embedding.
2 Background and Related Work
@@
This study delves into the problem of species-aware DNA embedding, aiming to create a model that maps each DNA sequence as a fixed-size numerical vector in a vector space, where sequences from distinct species are naturally clustered and segregated. A DNA sequence is essentially a string composed of four unique characters: A, T, C, and G.
@@
Existing works highly rely on descriptive textual features [Kang et al., 2015, 2019; Nissen et al., 2021] and pre-trained K-mer embeddings [Ng, 2017; Ren et al., 2022; Han et al., 2022] to compute DNA embeddings. A representative descriptive textual feature is Tetra-Nucleotide Frequency (TNF), a $256$-dimensional vector where each position represents the frequency of each unique $4$-mer (e.g., TTCA, AACG) in the input DNA sequence. Despite its simplicity and effectiveness, this method is limited since it is not trainable to better fit downstream applications. Besides, our empirical analysis also suggests that a naive trainable model based on TNF, such as a Variational AutoEnoder [Kingma and Welling, 2013] with TNF as input, results in worse embeddings compared to TNF. With the success of Word2Vec [Mikolov et al., 2013], pre-trained Kmer embeddings have gained popularity in computing DNA embeddings for various applications [Ng, 2017; Ren et al., 2022; Han et al., 2022]. However, the emergence of deep learning advancements such as ELMo and BERT [Peters et al., 2018; Devlin et al., 2018] highlights the limitations of static word embeddings compared to contextual embeddings produced by foundation models. Recently, Genome Foundation Models such as DNABERT-2 and HyenaDNA have demonstrated their prowess in genome analysis [Ji et al., 2021; Dalla-Torre et al., 2023; Nguyen et al., 2023; Zhou et al., 2023]. However, in the absence of task-specific data, these models struggle, largely due to the mismatch between their language-modeling training objectives and the goal of producing distinctive embeddings [Li et al., 2020].
In response, we turn to contrastive learning [Reimers and Gurevych, 2019; Gao et al., 2021; Chen et al., 2020; Lee et al., 2020] with genome foundation model and introduce the curriculum contrastive learning (C2LR) strategy with the Manifold Instance Mixup (MI-Mix) training objective.
@@ 3 Model
The proposed Curriculum Contrastive Learning ($\text{C}^{2}$LR) splits the training process into two phases, gradually creating more challenging anchors. In phase I, we apply an effective contrastive learning method named Weighted SimCLR based on SimCLR and Hard-Negative sampling strategy (Sec. 3.1). In phase II, we propose the Manifold Instance Mixup method which creates more challenging anchors by mixing intermediate hidden states of inputs in a randomly selected hidden layer of the model (Sec. 3.2). Implementation details of DNABERT-S are presented in Sec. 3.3.
Notation: Let $x_{i}$ define an input sample. Consider a batch $\{(x_{i},x_{i^{+}})\}_{i=1}^{B}$, where $B$ is the batch size and $(x_{i},x_{i^{+}})$ represents a pair of samples that we consider to be similar (a.k.a., positive pair). In our setting, a positive pair $(x_{i},x_{i^{+}})$ represents two non-overlapping DNA sequences from the same genome. Let $f(\cdot)$ define the embedding model, which takes $x_{i}$ as input and computes fixed-size embedding $f(x_{i})$.
@@ 3.1 Weighted SimCLR
Figure 2: Overview of DNABERT-S’s training process. We construct training data from massive reference genomes and train DNABERT-S with the proposed Curriculum Contrastive Learning (C2LR) strategy that progressively provides more challenging contrastive anchors to the model in two different phases. We propose the Manifold Instance Mixup (MI-Mix) objective that mixes the intermediate hidden states of different inputs to construct the mixed and challenging contrastive anchor.
SimCLR [Chen et al., 2020] is a simple and effective framework for contrastive learning. For an anchor $x_{i}$ in batch $\{(x_{i},x_{i^{+}})\}_{i=1}^{B}$, SimCLR treats all the other 2B22B-2 samples in the same batch as negative samples. It encourages the model to increase the anchor’s similarity with its positive sample $x_{i^{+}}$ and reduces its similarity with the negative samples. It treats all negative samples equally. However, recent works [Zhang et al., 2021] have suggested that hard negatives that are closer to the anchor in the representation space offer more informative learning contrasts. Therefore, Weighted SimCLR [Zhang et al., 2021] gives higher weights to negative samples that are closer to the anchor. To align with subsequent sections, we introduce the virtual labels. The label for $(x_{i},x_{i^{+}})$ is $v_{i}\in\{0,1\}^{B}$, where $v_{i,i}=1$ indicates positive samples, and $v_{i,j\neq i}=0$ indicates negative samples. The Weighted SimCLR loss for $x_{i}$ is defined as:
[…]
where $\tau$ denotes the temperature and $\operatorname{s}(\cdot,\cdot)$ denotes the cosine similarity between two inputs. Weights $\alpha_{ij}$ denotes the relative importance of $x_{j}$ for optimizing the contrastive loss of anchor $x_{i}$ among all the 2B22B-2 negative samples. A negative sample that is closer to the anchor receives a higher weight. We set $\alpha_{ii^{+}}=1$ and compute $\alpha_{ij}$ as:
[…]
@@
[…]
3.2 Curriculum Contrastive Learning ($\text{C}^{2}$LR)
In this part, we introduce our curriculum contrastive learning ($\text{C}^{2}$LR) method. Curriculum learning is an effective training method that first presents easy training batches and then progresses to more challenging ones [Hacohen and Weinshall, 2019]. Recent studies have successfully applied this technique to both positive pairs [Ye et al., 2021; Roy and Etemad, 2023] and negative pairs [Chu et al., 2021] in contrastive learning. We take this approach a step further by applying it to contrastive anchors, effectively using it for both types of pairs at the same time.
As shown in Figure 2, our $\text{C}^{2}$LR method includes two training phases, with anchors becoming progressively more challenging. In phase I, we use the Weighted SimCLR introduced in Sec. 3.1. In phase II, we propose the Manifold Instance Mixup (MI-Mix) method to mix up anchor instances in a random hidden layer, motivated by the instance mixup (i-Mix) method [Lee et al., 2020].
The i-Mix method mixes anchors at the input layer to create more challenging positive and negative pairs. It only uses the samples from $\{x_{i}\}_{i=1}^{B}$ as anchors and only considers the positive and negative samples from $\{x_{i^{+}}\}_{i=1}^{B}$. Otherwise, it nearly doubles the memory or training time compared to the Weighted SimCLR method in Sec. 3.1 (see Appendix A for details). To perform mixup within the anchor space, i-Mix first shuffles $\{(x_{i},v_{i})\}_{i=1}^{B}$ to generate $\{(\hat{x}_{i},\hat{v}_{i})\}_{i=1}^{B}$, i.e. a random permutation of $\{(x_{i},v_{i})\}_{i=1}^{B}$. Then for each anchor $(x_{i},v_{i})$, i-Mix mixes it with $(\hat{x}_{i},\hat{v}_{i})$ through weighted sum. The mixing weight $\lambda_{i}$ is drawn from $\text{Beta}(\alpha,\alpha)$, where $\alpha$ is a hyperparameter.

### v3

# DNABERT-S: Pioneering Species Differentiation with Species-Aware DNA Embeddings

@@ Abstract
We introduce DNABERT-S, a tailored genome model that develops species-aware embeddings to naturally cluster and segregate DNA sequences of different species in the embedding space. Differentiating species from genomic sequences (i.e., DNA and RNA) is vital yet challenging, since many real-world species remain uncharacterized, lacking known genomes for reference. Embedding-based methods are therefore used to differentiate species in an unsupervised manner. DNABERT-S builds upon a pre-trained genome foundation model named DNABERT-2. To encourage effective embeddings to error-prone long-read DNA sequences, we introduce Manifold Instance Mixup (MI-Mix), a contrastive objective that mixes the hidden representations of DNA sequences at randomly selected layers and trains the model to recognize and differentiate these mixed proportions at the output layer. We further enhance it with the proposed Curriculum Contrastive Learning (C2LR) strategy. Empirical results on 23 diverse datasets show DNABERT-S’s effectiveness, especially in realistic label-scarce scenarios. For example, it identifies twice more species from a mixture of unlabeled genomic sequences, doubles the Adjusted Rand Index (ARI) in species clustering, and outperforms the top baseline’s performance in 10-shot species classification with just a 2-shot training. Model, codes, and data is publicly available at https://github.com/MAGICS-LAB/DNABERT_S.
@@ 1 Introduction
Accurate differentiation of species from genomic sequences is a critical task in biology and ecology, supporting efforts in biodiversity conservation, epidemiology, understanding evolutionary processes, and exploring the roles of microbiomes in health and disease. Traditional methods for species identification rely heavily on well-characterized reference genomes for comparative analysis. Thus, they are limited due to the vast and largely unexplored genetic diversity present in natural environments.
A prime example is metagenomics binning. Metagenomics binning [Kang et al., 2015, 2019; Nissen et al., 2021; Meyer et al., 2022; Lamurias et al., 2023] is a pivotal process in microbiome research, aiming to group DNA sequences by species from complex mixtures containing DNA from potentially thousands of distinct, often uncharacterized species. In this context, effective DNA embeddings that can accurately segregate and cluster DNA sequences are more suitable than the methods that rely on known reference genomes for comparison and alignment.
@@
Despite the critical role of DNA embeddings in various scenarios, there is a notable deficiency in the development of effective methods. Current approaches to achieving DNA embeddings include: 1) Descriptive textual features [Kang et al., 2015, 2019; Nissen et al., 2021], 2) Pre-trained Kmer embeddings [Ng, 2017; Ren et al., 2022; Han et al., 2022], and 3) Genome foundation models [Ji et al., 2021; Nguyen et al., 2023; Zhou et al., 2023]. The first two methods, while straightforward, often fail to capture complex semantic relationships inherent in genomic data. Genome foundation models [Ji et al., 2021; Nguyen et al., 2023; Zhou et al., 2023], despite their successes in various genomic tasks through model fine-tuning, generally fail to develop embeddings that can discriminate certain properties like species. This is largely due to a mismatch between their pre-training objectives [Radford et al., 2019; Devlin et al., 2018] and the specific application scenarios. Our empirical analysis, as detailed in Table 1, reveals that in many scenarios, existing genome foundation models even underperform simple textual features.
@@
In this work, we introduce DNABERT-S, a specialized genome model that harnesses the capabilities of genome foundation models to generate species-aware DNA embeddings. As depicted in Figure 1, DNABERT-S distinguishes itself from other methods by its ability to effectively cluster and separate different species within the embedding space. This enhanced performance stems from the proposed Manifold Instance Mixup (MI-Mix) loss and Curriculum Contrastive Learning (C2LR) strategy. Contrastive learning enables the model to discern between similar and dissimilar DNA sequences, and curriculum learning incrementally presents more challenging training samples, fostering better learning and generalization. The training of DNABERT-S includes two phases. In the first phase, we adopt a Weighted SimCLR Chen et al. [2020]; Zhou et al. [2022] training objective to encourage the model to group DNA sequences from the same species and separate DNA sequences from distinct species. In the second phase, we introduce Manifold Instance Mixup (MI-Mix), which mixes anchor instances at a randomly selected layer to create more challenging anchors for contrastive training.
@@
To rigorously evaluate the models, we assemble a comprehensive benchmark that includes thousands of species, capturing the diversity of natural microbial communities. This benchmark incorporates complex samples from CAMI2 [Meyer et al., 2022], a leading metagenomics binning benchmark, and extensive reference genomes from Genbank [Benson et al., 2012]. Thus, it includes both natural yet error-prone long-read sequences and well-curated yet potentially biased reference genomes.
@@
We evaluate the embedding quality from entirely unsupervised problems to classification tasks with abundant labels. Experimental results indicate the effective performance of DNABERT-S. Compared to the strongest existing method, DNABERT-S doubles its performance in the clustering task and achieves better performance with only $20\%$ of labeled data in the classification task (e.g., 2-shot v.s. 10-shot). Notably, in metagenomics binning, DNABERT-S is able to recover over $40\%$ and $80\%$ of species with an F1 score of over $0.5$ respectively from synthetics and more realistic datasets, which is also one time more than the strongest baseline. We also show that a simple K-Nearest-Neighbors species classifier, which is trained on DNABERT-S embeddings of a small portion of each target genome, can slightly outperform a well-established traditional method MMseqs2 [Steinegger and Söding, 2017] that relies on the entire reference genomes of target species for classification.
@@
Our contribution can be summarized as follows: 1) We demonstrate the effectiveness of genome foundation models in learning DNA embeddings, opening new avenues for various genomics research problems; 2) We introduce DNABERT-S, a model that develops distinctly better embeddings for species differentiation; 3) We propose the Curriculum Contrastive Learning (C2LR) strategy with the Manifold Instance Mixup (MI-Mix) loss; 4) We publish a large-scale evaluation benchmark.
2 Background
@@
This study aims to build a species-aware DNA embedding model that maps each DNA sequence as a fixed-size numerical vector in an embedding space, where sequences from distinct species are naturally clustered and segregated. A DNA sequence is essentially a sentence composed of four unique characters: A, T, C, and G.
@@
Existing works highly rely on descriptive textual features [Kang et al., 2015, 2019; Nissen et al., 2021] and pre-trained K-mer embeddings [Ng, 2017; Ren et al., 2022; Han et al., 2022] to compute DNA embeddings. A representative descriptive textual feature is Tetra-Nucleotide Frequency (TNF), a $256$-dimensional vector where each position represents the frequency of each unique $4$-mer (e.g., TTCA, AACG) in the input DNA sequence. Despite its simplicity and effectiveness, this method is limited since it singly relies on the $4$-mer frequency and is not trainable to better fit downstream applications. Besides, our empirical analysis also suggests that a naive trainable model based on TNF, such as a Variational AutoEnoder [Kingma and Welling, 2013] with TNF as input, results in worse embeddings compared to TNF. With the success of Word2Vec [Mikolov et al., 2013], pre-trained Kmer embeddings have gained popularity in computing DNA embeddings for various applications [Ng, 2017; Ren et al., 2022; Han et al., 2022]. However, the emergence of deep learning advancements such as ELMo and BERT [Peters et al., 2018; Devlin et al., 2018] highlights the limitations of static word embeddings compared to contextual embeddings produced by foundation models.
Recently, genome foundation models such as DNABERT-2 and HyenaDNA have demonstrated their effectiveness in genome analysis [Ji et al., 2021; Dalla-Torre et al., 2023; Nguyen et al., 2023; Zhou et al., 2023]. However, these models do not naturally develop discriminative embeddings, largely due to the discrepancy between their language-modeling training objectives and the goal of segregating sequences in the embedding space [Li et al., 2020]. To leverage the power and potential of genome foundation models, we tailor a contrastive learning method [Reimers and Gurevych, 2019; Gao et al., 2021; Chen et al., 2020; Lee et al., 2020] for DNA embedding learning by introducing the curriculum contrastive learning (C2LR) strategy with the Manifold Instance Mixup (MI-Mix) training objective.
@@ 3 Model
The proposed Curriculum Contrastive Learning ($\text{C}^{2}$) splits the training process into two phases, gradually creating more challenging anchors. In phase I, we apply an effective contrastive learning method named Weighted SimCLR based on SimCLR and Hard-Negative sampling strategy (Sec. 3.1). In phase II, we propose the Manifold Instance Mixup method which creates more challenging anchors by mixing intermediate hidden states of inputs in a randomly selected hidden layer of the model (Sec. 3.2). Implementation details of DNABERT-S are presented in Sec. 3.3.
Notation: Let $x_{i}$ be an input sample. Given a batch $\{(x_{i},x_{i^{+}})\}_{i=1}^{B}$, where $B$ is the batch size and $(x_{i},x_{i^{+}})$ represents a pair of similar samples (a.k.a., positive pair). In our setting, a positive pair $(x_{i},x_{i^{+}})$ represents two non-overlapping DNA sequences from the same genome. Let $f(\cdot)$ define the embedding model, which takes $x_{i}$ as input and computes fixed-size embedding $f(x_{i})$.
@@ 3.1 Weighted SimCLR
SimCLR [Chen et al., 2020] is a simple and effective framework for contrastive learning. For an anchor $x_{i}$ in batch $\{(x_{i},x_{i^{+}})\}_{i=1}^{B}$, SimCLR treats all the other $22B-$ 2 samples in the same batch as negative samples. It encourages the model to increase the anchor’s similarity with its positive sample $x_{i^{+}}$ and reduces its similarity with the negative samples. It treats all negative samples equally. However, recent works [Zhang et al., 2021] have suggested that hard negatives that are closer to the anchor in the representation space offer more informative learning contrasts. Therefore, Weighted SimCLR [Zhang et al., 2021] gives higher weights to negative samples that are closer to the anchor.
Figure 2: Overview of DNABERT-S’s training process. We construct training data from massive reference genomes and train DNABERT-S with the proposed Curriculum Contrastive Learning (C2LR) strategy that progressively provides more challenging contrastive anchors to the model in two different phases. We propose the Manifold Instance Mixup (MI-Mix) objective that mixes the intermediate hidden states of different inputs to construct more challenging contrastive anchor.
To align with subsequent sections, we introduce the virtual labels. The label for $(x_{i},x_{i^{+}})$ is $v_{i}\in\{0,1\}^{B}$, where $v_{i,i}=1$ indicates positive samples, and $v_{i,j\neq i}=0$ indicates negative samples. The Weighted SimCLR loss for $x_{i}$ is defined as:
$\ell(f(x_{i}),v_{i})=-\sum_{n=1}^{B}v_{i,n}\log\frac{\exp\left(\operatorname{s}\left(f(x_{i}),f(x_{n^{+}})\right)/\tau\right)}{\sum_{j\neq i}\alpha_{ij}\exp\left(\operatorname{s}\left(f(x_{i}),f(x_{j})\right)/\tau\right)},$ (1)
where $\tau$ denotes the temperature and $\operatorname{s}(\cdot,\cdot)$ denotes the cosine similarity between two inputs. Weights $\alpha_{ij}$ denotes the relative importance of $x_{j}$ for optimizing the contrastive loss of anchor $x_{i}$ among all the $22B-$ 2 negative samples. A negative sample that is closer to the anchor receives a higher weight. We set $\alpha_{ii^{+}}=1$ and compute $\alpha_{ij}$ as:
$\alpha_{ij}=\frac{\exp\left(\operatorname{s}\left(f(x_{i}),f(x_{j})\right)/\tau\right)}{\frac{1}{2B-2}\sum_{k\neq i,i^{+}}\exp\left(\operatorname{s}\left(f(x_{i}),f(x_{k})\right)/\tau\right)}.$
@@
$\mathcal{L}=\frac{1}{2B}\sum_{i=1}^{B}(\ell(f(x_{i}),v_{i})+\ell(f(x_{i^{+}}),v_{i})).$ (2)
3.2 Curriculum Contrastive Learning ($\text{C}^{2}$)
In this part, we introduce our curriculum contrastive learning ($\text{C}^{2}$) method. Curriculum learning is an effective training method that first presents easy training batches and then progresses to more challenging ones [Hacohen and Weinshall, 2019]. Recent studies have successfully applied this technique to both positive pairs [Ye et al., 2021; Roy and Etemad, 2023] and negative pairs [Chu et al., 2021] in contrastive learning. We take this approach a step further by applying it to contrastive anchors, effectively using it for both types of pairs at the same time.
As shown in Figure 2, our $\text{C}^{2}$ method includes two training phases, with anchors becoming progressively more challenging. In phase I, we use the Weighted SimCLR introduced in Sec. 3.1. In phase II, we propose the Manifold Instance Mixup (MI-Mix) method to mix up anchor instances in a random hidden layer, motivated by the instance mixup (i-Mix) method [Lee et al., 2020].
The i-Mix method mixes anchors at the input layer to create more challenging positive and negative pairs. It only uses the samples from $\{x_{i}\}_{i=1}^{B}$ as anchors and only considers the positive and negative samples from $\{x_{i^{+}}\}_{i=1}^{B}$. Otherwise, it nearly doubles the memory or training time compared to the Weighted SimCLR method in Sec. 3.1 (see Appendix D for details). To perform mixup within the anchor space, i-Mix first shuffles $\{(x_{i},v_{i})\}_{i=1}^{B}$ to generate $\{(\hat{x}_{i},\hat{v}_{i})\}_{i=1}^{B}$, i.e. a random permutation of $\{(x_{i},v_{i})\}_{i=1}^{B}$. Then for each anchor $(x_{i},v_{i})$, i-Mix mixes it with $(\hat{x}_{i},\hat{v}_{i})$ through weighted sum. The mixing weight $\lambda_{i}$ is drawn from $\text{Beta}(\alpha,\alpha)$, where $\alpha$ is a hyperparameter.